# Chern character of Schur bundles

### v1

@@ 3.1. Chern character of symmetric and exterior powers
$h=h(\nu)=\nu_{1}+\cdots\nu_{\ell}$
to be the height of the Young diagram $\mathbf{a}_{\nu}$, i.e. the number of rows. If $\mathbf{b}=(b_{1},\dots,b_{h})$ is $h$-tuple of integers, we put
@@
$\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{m}E)=\sum_{\nu\vdash n}\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}\,\operatorname{ch}^{(\nu)}(E)$ (3.3)
@@
$\operatorname{ch}_{n}\left(\bigwedge^{m}E\right)=\sum_{\nu\vdash n}\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|-\nu_{1}}\,\operatorname{ch}^{(\nu)}(E),$ (3.4)
where the sums run over all $\mathbf{b}=(b_{1},\dots,b_{h(\nu)})$ such that $0<b_{i}\leq a_{i}$ for every $i$, and $|\mathbf{b}|=\sum_{i}b_{i}$.
@@
We can be more precise about the bounds $0<b_{i}\leq a_{i}$. In fact, if $a_{i}>1$, then we have $0<b_{i}\leq a_{i}-1$, as it will be clear from the proof. The extremal case $b_{i}=a_{i}$ only appears when $a_{i}=1$, i.e. $a_{i}-1=0$, see also Remark 3.7.
In the proof we will need the following well-known binomial identities.
@@
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R+N-i}{R}=\binom{R+N+A+1}{R+A+1}$ (3.5)
and
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R}{N-i}=\binom{R-A-1}{N}$. (3.6)
Proof of Proposition 3.8 (symmetric power case).
It is enough to prove that
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}_{\nu}}(\mathsf{\mathbf{s}}_{m}))=\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}$. (3.7)
We proceed by induction on $h=h(\nu)$. As base of the induction, let us prove the $\mathbf{a}_{\nu}=(a)$. By definition and (3.2),
@@
For the inductive step, let us first assume that there exists an index $i$ such that $a_{i}>1$. Without loos of generality, we can assume that it is $a_{1}>1$. In fact, since $\mathsf{D}_{a_{i}}$ and $\mathsf{D}_{a_{j}}$ commute we can re-order the indices. Let us then write $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$, and let us set $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$. By definition we have
@@
Proof of Proposition 3.8 (exterior power case).
The proof is analogous to the previous one, so we will only sketch it.
First of all, let us deal with the case $\mathbf{a}_{\nu}=(a)$. As before, we have different behavior if $a=1$ or $a>1$. Since
$\operatorname{rk}\mathsf{D}_{a}\left(\bigwedge^{m}E\right)=\sum_{k01}^{m}(-1)^{k+1}k^{a-1}\binom{r}{m-k}=\left\{\begin{array}[]{ll}\binom{r-1}{m-1}&\mbox{if }a=1\\ \sum_{b=1}^{a-1}\mathsf{\mathbf{E}}(a-1,b-1)\binom{r-a}{m-a-b}&\mbox{if }a>1\,.\end{array}\right$.
Let now $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$ and assume that for at least one index $i$ we have $a_{i}>1$. Without loss of generality, it is $a_{1}>1$. If $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$, then
$\operatorname{rk}\mathsf{D}_{\mathbf{a}_{\nu}}\left(\bigwedge^{m}E\right)=$ $\,\sum_{k=1}^{m}(-1)^{k+1}k^{a_{1}-1}\operatorname{rk}\mathsf{D}_{\mathbf{a}^{\prime}}\left(\bigwedge^{m-k}E\right)$
@@
where the last equality follows from (3.6).
@@
$\operatorname{rk}\mathsf{D}_{1}^{h}\left(\bigwedge^{m}E\right)=\binom{r-1}{r-h}$,
@@
Remark 3.11.
In [Svr93], Svrtan obtains a similar formulas involving Stirling numbers of the second kind. It is known that Eulerian numbers and Stirling numbers of the second kind are related, so our formula could have deduced from his. Nevertheless, notice that our proof is completely different from his.
4. Twisted Chern character of Schur bundles
Let $\mathsf{C}^{E}$ be the characteristic class defined in (3.1). We can twist it in the following way,
$\mathsf{C}^{\mathbb{P}E}\colon\Lambda\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto k(\mathsf{\mathbf{S}}^{\alpha}E):=\exp\left(-\frac{|\alpha|}{r}\operatorname{ch}_{1}(E)\right)\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E).$
It is easy to check that this is again a characteristic class. Notice that in this case $\mathsf{C}^{\mathbb{P}E}$ is only same-degree additive. The name and the notation are motivated by [Mar20], where these classes are used as Chern classes of projective bundles. More precisely, for any (twisted) vector bundle $F$, Markman defines
Our definition is equivalent to the above one thanks to Corollary 3.3. The first three values of $k(F)$ in terms of Chern classes are
and
Notice that $2\operatorname{rk}(F)\,k_{2}(F)=\Delta(F)$ is the discriminant of $F$. As a direct application of Corollary 1.4, we get the following formula for the class $k$ of Schur bundles.
Theorem 4.1.
With notations as above,
Example 4.2.
so that, for instance,
$k_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\,k_{2}(E)\,;$
$k_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\,k_{3}(E)\,;$
$k_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\,k_{4}(E)+\frac{1}{2}k_{2}^{2}(E)\,.$
Remark 4.3.
from which it follows, for example, that
In [DFO25a] we proved that the classes $\mathsf{L}_{2}(F)$ and $\mathsf{L}_{3}(F)$ are eigen-classes with respect to taking the Schur functors of a bundle. Moreover, we also remarked that the same did not hold for higher-degree classes and speculated that similar formulas would work for the projection of the classes $\mathsf{L}_{n}(\mathsf{\mathbf{S}}^{\alpha}E)$ onto the $\mathsf{L}_{n}(E)$-component.

### v2

@@ 4. Chern characters of symmetric and exterior powers
$h=h(\nu)=\nu_{1}+\cdots+\nu_{\ell}$
to be the height of the Young diagram $\mathbf{a}_{\nu}$, i.e. the number of rows. If $\mathbf{b}=(b_{1},\dots,b_{h})$ is an $h$-tuple of integers, we put
@@
Similarly, we write
$k\leq\mathbf{b}\leq\mathbf{c}\implies k\leq b_{i}\leq c_{i}\quad\forall i=1,\dots,h$
and
$|\mathbf{b}|=b_{1}+\cdots+b_{h}.$
Let us state the formulas. We start with the symmetric power.
@@
$\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{m}E)=\sum_{\nu\vdash n}\sum_{1\leq\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}\,\operatorname{ch}^{(\nu)}(E)$
@@
$k_{n}(\mathsf{\mathbf{S}}^{m}E)=\sum_{\begin{subarray}{c}\nu\vdash n\\ \nu_{1}=0\end{subarray}}\sum_{1\leq\mathbf{b}\leq\overleftarrow{\mathbf{a}_{\nu}}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-1}{r+|\mathbf{a}_{\nu}|-1}\,k^{(\nu)}(E).$
For exterior products, we have instead the following statement.
Proposition 4.3.
With notations as above, we have
$\operatorname{ch}_{n}\left(\mathop{\bigwedge\nolimits^{\!m}}E\right)=\sum_{\nu\vdash n}\sum_{1\leq\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|-\nu_{1}}\,\operatorname{ch}^{(\nu)}(E)$
and
$k_{n}\left(\mathop{\bigwedge\nolimits^{\!m}}E\right)=\sum_{\begin{subarray}{c}\nu\vdash n\\ \nu_{1}=0\end{subarray}}\sum_{1\leq\mathbf{b}\leq\overleftarrow{\mathbf{a}_{\nu}}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|}\,k^{(\nu)}(E).$
@@
We can be more precise about the bound $\mathbf{b}\leq\mathbf{a}_{\nu}$. In fact, if $a_{i}>1$, then we have $b_{i}\leq a_{i}-1$, as it will be clear from the proof. The extremal case $b_{i}=a_{i}$ only appears when $a_{i}=1$, i.e. $a_{i}-1=0$, see also Remark 4.1. This also explains why in the formula for the twisted Chern character we can restrict to $\mathbf{b}\leq\overleftarrow{\mathbf{a}_{\nu}}$.
4.1. Proof of Proposition 4.2
We will need the following well-known binomial identity.
@@
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R+N-i}{R}=\binom{R+N+A+1}{R+A+1}$.
First of all, it is enough to prove that
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}_{\nu}}(\mathsf{\mathbf{s}}_{m}))=\sum_{1\leq\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}$. (4.2)
In fact, this is exactly the coefficient in front of $\operatorname{ch}^{(\nu)}(E)$ in the formula for the Chern character, while for the twisted Chern character it is enough to notice that $k_{1}(E)=0$, so that every partition $\nu$ such that $\nu_{1}\geq 1$ does not contribute to the formula. (The shift $\overleftarrow{\mathbf{a}_{\nu}}$ in the formula for the twisted Chern character is explained in Remark 4.4.)
We proceed by induction on $h=h(\nu)$. As base of the induction, let us prove the case $\mathbf{a}_{\nu}=(a)$. By definition and (4.1),
@@
For the inductive step, let us first assume that there exists an index $i$ such that $a_{i}>1$. Without loss of generality, we can assume that it is $a_{1}>1$. In fact, since $\mathsf{D}_{a_{i}}$ and $\mathsf{D}_{a_{j}}$ commute we can re-order the indices. Let us then write $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$, and let us set $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$. By definition we have
@@
4.2. Proof of Proposition 4.3
We will need the following well-known binomial identity.
Lemma 4.6.
For any non-negative integers $A$, $R$ and $N$,
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R}{N-i}=\binom{R-A-1}{N}$.
The proof goes as in § 4.1, so we will only comment on the due changes. First of all, as before it is enough to prove that
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}_{\nu}}(\mathsf{\mathbf{s}}_{1^{m}}))=\sum_{1\leq\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|}$. (4.3)
This already implies the formula for the Chern character, while the formula for the twisted Chern character is obtained by noticing that $k_{1}(E)=0$ as before.
Let us start by settling the base of the induction and prove the case $\mathbf{a}_{\nu}=(a)$. As before, we have different behaviour if $a=1$ or $a>1$:
$\operatorname{rk}\mathsf{D}_{a}\left(\mathop{\bigwedge\nolimits^{\!m}}E\right)=\sum_{k=1}^{m}(-1)^{k+1}k^{a-1}\binom{r}{m-k}=\left\{\begin{array}[]{ll}\binom{r-1}{m-1}&\mbox{if }a=1\\ \sum_{b=1}^{a-1}\mathsf{\mathbf{E}}(a-1,b-1)\binom{r-a}{m-a-b}&\mbox{if }a>1\end{array}\right$.
where we used Lemma 4.6.
Put now $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$ and assume that for at least one index $i$ we have $a_{i}>1$. Without loss of generality, it is $a_{1}>1$. If $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$, then
$\operatorname{rk}\mathsf{D}_{\mathbf{a}_{\nu}}\left(\mathop{\bigwedge\nolimits^{\!m}}E\right)=$ $\,\sum_{k=1}^{m}(-1)^{k+1}k^{a_{1}-1}\operatorname{rk}\mathsf{D}_{\mathbf{a}^{\prime}}\left(\mathop{\bigwedge\nolimits^{\!m-k}}E\right)$
@@
where the last equality follows from Lemma 4.6.
@@
$\operatorname{rk}\mathsf{D}_{1}^{h}\left(\mathop{\bigwedge\nolimits^{\!m}}E\right)=\binom{r-1}{r-h}$,